A car starting from rest accelerates uniformly and attains a speed of 80m/s in 30s. It maintains this steady speed for another 30s. Then slows down uniformly until it comes to rest in the next 40s. What is the total distance covered?

1
Expert's answer
2022-01-18T09:36:57-0500

"d_1=\dfrac {vt_1}2=\dfrac {80·30}2=1200\text{ m}.\\\space\\
d_2=vt_2=80·30=2400\text{ m}.\\\space\\
d_3=\dfrac{vt_3}{2}=\dfrac{80·40}{2}=1600\text{ m}.\\\space\\
D=d_1+d_2+d_3=5200\text{ m}."